sam has a bag with 2 red, 7 yellow, and 5 blue marbles in it. if he pulls out 3 marbles, what is the probability that at least one of them is blue?
5/14 x 5/13 x 5/12 = 125/2184

Found 2 solutions by stanbon, scott8148:
Answer by stanbon(75887)

sam has a bag with 2 red, 7 yellow, and 5 blue marbles in it. if he pulls out 3 marbles, what is the probability that at least one of them is blue?
-------------------
P(at least one is blue) = 1 - P(none of the 3 is blue)
-----
= 1 - 9C3/14C3
----
= 1 - 84/364
---
= 1 - 0.2308
===============
= 0.7692
=================
Cheers,
Stan H.







5/14 x 5/13 x 5/12 = 125/2184

Answer by scott8148(6628)

number of ways to draw 3 marbles ___ 14C3 = 364

number of ways to draw 3 marbles, with no blue ___ 9C3 = 84

so 280 draws (364 - 84) have at least one blue

probability of at least one blue ___ 280 / 364 = .77 (approx)